OpenStudy (anonymous):

Differential help (D^2-1)y = 10sin(3x)

OpenStudy (anonymous):

im trying to find my Yp

OpenStudy (anonymous):

I struggle a bit with this notation, not really used to Euler: \[\Large y''-y=10\sin(3x) \] ?

OpenStudy (anonymous):

ive set up my equation to be -90Acos(3x)-90Bsin(3x)=10sin(3x) I just need to solve this for A and B

OpenStudy (anonymous):

im not sure what to set my equilibrium equations equal to

OpenStudy (anonymous):

What did you take as a guess?

OpenStudy (anonymous):

\[\Large A\sin(3x)+B\cos(3x) \]?

OpenStudy (anonymous):

Because that would change your differentiation a bit, you should end up at:\[\Large -10A\sin(3x)-10B\cos(3x)=10\sin(3x) \]

OpenStudy (anonymous):

but you then have to take the derivative of this twice and when you plug that into D^2-1 that should end you up where i was at

OpenStudy (anonymous):

Well, as I requested in my opening post; I really don't see the Euler Notation too often, but if I understand this problem correctly then it's asking for: \[\Large y''-y=10\sin(3x) \] is this correct?

OpenStudy (anonymous):

Or did I misinterpret the Euler Notation? I think I remember that \[\Large D^2f=f'' \]

OpenStudy (unklerhaukus):

what did you get for \(y_h\) ?

OpenStudy (anonymous):

i dont even know what yh is

OpenStudy (anonymous):

yc = ae^x+be^(-x)

OpenStudy (unklerhaukus):

yeah y_h is the same as y_c

OpenStudy (anonymous):

ok that makes sense so how do i solve for my A and B values?

OpenStudy (anonymous):

\[\Large y''-y=0 \\ \Large \lambda^2-1=0 \] So I agree But your guess above still seems wrong to me. \[\Large y_p=A\sin(3x)+B\cos(3x) \] Differetiate that twice, I don't see how your 90 shows up.

OpenStudy (anonymous):

im not sure how you came up with this but go ahead and continue

OpenStudy (anonymous):

\[\Large y_p'=3A\cos(3x)-3B\sin(3x) \\ \Large y_p''=-9A\sin(3x)-9B\cos(3x) \] Such that: \[\Large y_p''-y_p=-10A\sin(3x)-10B\cos(3x)=10\sin(3x) \]

OpenStudy (anonymous):

B=0, A=-1

OpenStudy (anonymous):

wait can you go over the solving part

OpenStudy (anonymous):

can you tell me where you've problems? At the guess? At the differentiation? Or at the last step where you try to figure out the constants.

OpenStudy (anonymous):

the last step. solving for the A and B variable.

OpenStudy (anonymous):

Well basically all you do is set things equal: \[\Large -10A\sin(3x)-10B\cos(3x)=10\sin(3x) \] Remember that the equal sign means the LHS (Left hand side) must be equal to the RHS. So on the right hand side we don't have a cos(3x) term, how can we make it disappear from the left hand side? By settign B=0, Now we're left with: \[\Large -10A\sin(3x)=10\sin(3x) \] Solving for A gives A=-1

OpenStudy (anonymous):

so can just assume b to be 0? you dont have to prove this?

OpenStudy (anonymous):

Have you done partial fraction decomposition before?

OpenStudy (anonymous):

There you do basically the exact same thing, the idea behind this kind of differential equations it that you first solve the homogenous DE, which is usually easy to solve. Then you make a guess for the RHS. This step is necessary because a differential equation is defined through the differentials, so in anyway, the function cannot just completely disappear. Your guess will always include unknown constants, and you're trying to solve for these. You do this by plugging in your guess into the original DE and then make your coefficients match the RHS of the equation. In general case, this is proof enough, it's a basic concept of mathematics to match coefficients. For further examples see: Laws of Vieta Integrating Factors